ADNAN MUHAMMED

Grade 12,

The molecule diameter of nitrogen is 3.5*10^-8 cm. calculate the mean free path at temprature 27 degree centigrate and p=1 atm please sir ,answer this question..

 
please sir ,answer this question..

Grade:12

1 Answers

Susmita
5 years ago
p=nkT
Or,n=p/kT
Where p=1atm=1.013×105 Pa ,T=27+273=300K,k=1.38×10-23Js.
so n=2.4×10^25
Mean free path
\lambda =\frac{1}{ \sqrt{2} \pi (\sigma)^2 n }
where diameter
\sigma =3.5*10^{-8}cm=3.5×10-10m
so lamda=7.5×10-8m